# Quantum Key Leasing for PKE and FHE with a Classical Lessor

### 03 Jan 2024 v3

@@ 10.2.1 Proof Outline
When we move on to using an approximate version of ${\sf TI}$, we will have to be more carefull: negligible errors will accumulate additively on the measurement outcomes $\rho_{\mathsf{Delete}}$ after every ${\sf TI}$-measurement, but will stay negligible throughout the algorithm with our choice of parameters (see Section 11 for details) by the properties of ATI in Section 5.2. By these guarantees, we obtain $\mathbf{x}$ with $1$ probability in the end.
@@ 11.2 Extraction Algorithm
  • Let $\mathbf{x}_{\ell,j}^{\prime}$ be register that stores the final guess for the $j$-th entry of $\mathbf{x}_{\ell}$.
@@
          Run ${\sf ATI}_{1/2+\gamma,{\mathcal{P}},{\mathcal{D}}(g_{\ell,j})}^{\epsilon,\delta}$ on $\rho_{\mathsf{Delete}}$ wih respect to $\mathsf{pk}$, for some inverse-polynomial $\epsilon=\gamma/100$.
@@ 11.3 Analysis of the Extractor
  • If ${\sf ATI}^{\epsilon,\delta}_{{\mathcal{P}},{\mathcal{D}}_{2.k},1/2+\gamma}\,\rho$ outputs 1 and a leftover state $\rho^{\prime}$, then $\operatorname{Tr}[{\sf ATI}^{\epsilon,\delta}_{{\mathcal{P}},{\mathcal{D}}(g_{\ell,j})),1/2+\gamma-3\epsilon}\,\rho^{\prime}]\geq 1-O(\eta_{0}^{\prime})$.
where $\eta_{0}^{\prime}=\eta_{0}^{3/2}\cdot{\sf poly}(\lambda)\cdot(1/\epsilon^{2}\cdot\ln(4/\delta))^{3}$ and $\eta_{0}$ is the statistical distances between $({\mathcal{D}}_{2.k},{\mathcal{D}}(g_{\ell,j}))$.
@@
This follows directly from 11.1, Lemma 5.8 and Theorem 5.11.
By Lemma 5.8, we have that if ${\sf ATI}^{\epsilon,\delta}_{{\mathcal{P}},{\mathcal{D}}_{2.k},1/2+\gamma}\,\rho$ outputs 1 and a leftover state $\rho^{\prime}$, then $\ \operatorname{Tr}[{\sf ATI}^{\epsilon,\delta}_{{\mathcal{P}},{\mathcal{D}}_{2.k},1/2+\gamma-3\epsilon}\,\rho^{\prime}]\geq 1-\delta$. Next, by Theorem 5.11, we have $\operatorname{Tr}[{\sf ATI}^{\epsilon,\delta}_{{\mathcal{P}},{\mathcal{D}}(g_{\ell,j})),1/2+\gamma}\,\rho^{\prime}]\geq\operatorname{Tr}[{\sf ATI}^{\epsilon,\delta}_{{\mathcal{P}},{\mathcal{D}}_{2.k},1/2+\gamma}\,\rho^{\prime}]-\eta_{0}^{\prime}$. Thus, overall we have $\operatorname{Tr}[{\sf ATI}^{\epsilon,\delta}_{{\mathcal{P}},{\mathcal{D}}(g_{\ell,j})),1/2+\gamma-3\epsilon}\,\rho^{\prime}]\geq 1-\delta-\eta_{0}^{\prime}=1-O(\eta_{0}^{\prime})$
@@
By Corollary 11.3, have that $\operatorname{Tr}[{\sf ATI}^{\epsilon,\delta}_{{\mathcal{P}},D(g_{i}),1/2+\gamma-4\epsilon}\,\rho]\geq 1-3\delta-O(\eta_{0}^{\prime})\geq 1-O(\eta_{0}^{\prime})$, where $\delta$ is exponentially small and $\eta_{0}^{\prime}={\sf poly}(\lambda)\cdot(\ln(4/\delta)/\epsilon^{2})^{3}\cdot(q^{-n}+B_{P}/B_{P^{\prime}}))$, according to 11.1.
@@
When the $L+1$-th measurement ${\sf ATI}^{\epsilon,\delta}_{{\mathcal{P}},D(g_{\ell,j}),1/2+\gamma}$ uses a correct $g_{\ell,j}$, then with probability $(1-L^{\prime}\cdot O(\eta_{1}\cdot\ln(4/\delta)/\epsilon))(1-O(\eta_{0}^{\prime}))$, we obtain outcome $1$, where $L^{\prime}$ is the number of times we get outcome 0 since the last time we get outcome 1, by the fact that $|\operatorname{Tr}({\mathcal{P}}\rho_{L-L^{\prime}})-\operatorname{Tr}({\mathcal{P}}\rho_{L})|\leq\lVert\rho_{L-L^{\prime}}-\rho_{L}\rVert_{\operatorname{Tr}}$ for all POVM measurements ${\mathcal{P}}$.
The total probability we extract all $k\cdot n$-entries of the LWE secret is: $(1-O(\eta_{0}^{\prime}))^{kn}\cdot(1-\delta)^{O(B_{X}\cdot kn)}\cdot(1-O(\eta_{1}\cdot\ln(4/\delta)/\epsilon))^{O(B_{X}\cdot kn)}$. Here, $(1-O(\eta_{0}^{\prime}))^{kn}$ is the total loss incurred by measurements ${\sf ATI}^{\epsilon,\delta}_{{\mathcal{P}},D(g_{\ell,j}),1/2+\gamma}$ with correct $g_{\ell,j}$’s, because we will make $kn$ (the dimension of the secret) number of such measurements; the error $(1-\delta)^{O(B_{X}\cdot kn)}\cdot(1-O(\eta_{1}\cdot\ln(4/\delta)/\epsilon))^{O(B_{X}\cdot kn)}$ is incurred by measurements ${\sf ATI}^{\epsilon,\delta}_{{\mathcal{P}},D(g_{\ell,j}),1/2+\gamma}$ with incorrect $g_{\ell,j}$’s, because we will make $O(B_{X}kn)$ (the dimension of the secret multiplied by the range of the secret) number of such measurements.
@@
Let us recall our parameter settings (the followings are all in terms of security parameter $\lambda$): $q^{-n}$ is exponentially small and thus $\eta_{1}$ is exponentially small; $B_{P}\cdot B_{X}^{4}/B_{P^{\prime}}$ is inverse superpolynomial by Section 9.1, $\epsilon=\frac{\gamma}{100B_{X}kn}$ and therefore the $\eta_{0}^{\prime}=O((\frac{B_{X}^{4}\cdot B_{P}}{B_{P^{\prime}}})^{\frac{3}{2}}\cdot{\sf poly}(\lambda))$ is inverse superpolynomial. $\delta$ is exponentially small. $k,n$ are polynomial. $B_{X}$ is quasipolynomial.
@@
$(1-O(\eta_{0}^{\prime}))^{kn}\cdot(1-\delta)^{O(B_{X}\cdot kn)}\cdot(1-O(\eta_{1}\cdot\ln(4/\delta)/\epsilon))^{O(B_{X}\cdot kn)}$
$\geq$ $(1-O(kn\cdot\eta_{0}^{\prime})\cdot(1-O(B_{X}kn\cdot\delta))\cdot(1-O(B_{X}kn\cdot\eta_{1}\cdot\ln(4/\delta)/\epsilon))$
$\geq$ $1-O(kn\cdot\eta_{0}^{\prime})$
@@ Appendix C Additional Missing Proofs
C.2 Proof for Theorem 5.11
We follow the proof in [Zha20] Theorem 6.5 with some modifications.
The difficulty is that ${\sf ATI}$ makes queries on a superposition of exponentially-many samples from the respective ${\mathcal{D}}_{b}$ distribution, whose indistinguishability does not follow from the indistinguishability of single samples. We tackle with this using the small range distribution technique in [Zha12].
$D_{0},D_{1}$ are distributions over ${\mathcal{R}}$, by distance $\eta$.
Hybrid 0
: This is the original world using ATI with respect to $D_{0}$ we denote the probability that our output is 1 as $\operatorname{Tr}[{\sf ATI}_{{\mathcal{P}},D_{0},\gamma}^{\epsilon,\delta}]$.
Hybrid 1
: we choose a random permutation $\Pi$ on ${\mathcal{R}}$. Let $D_{0}^{\Pi}(r)=D_{0}(\Pi(r))$. Since $D_{0}$ and $D_{0}^{\Pi}$ are identical distributions, the measurements ${\mathcal{P}}_{D_{0}}$ and ${\mathcal{P}}_{C^{\Pi}}$ are identical, and therefore so are their projective/threshold implementations. Thus, $\operatorname{Tr}[{\sf ATI}_{{\mathcal{P}},D_{0},\gamma}^{\epsilon,\delta}]$ and $\operatorname{Tr}[{\sf ATI}_{{\mathcal{P}},D^{\Pi}_{0},\gamma}^{\epsilon,\delta}]$ are identically distributed, for any $\gamma\in[0,1]$.
Hybrid 2
Now we change $\Pi$ to be the small-range functions $\Sigma=G\circ F$ of [Zha12], where $F:{\mathcal{R}}\to[s]$ and $G:[s]\to{\mathcal{R}}$ are random functions, and s is a parameter we will decide later.
By Theorem 6.6, Theorem 6.7 in [Zha20] (eprint version; originally in [Zha12, Yue13]), we have:
$\operatorname{Tr}[{\sf ATI}_{{\mathcal{P}},D_{0}^{\Sigma},\gamma}^{\epsilon,\delta}]\geq\operatorname{Tr}[{\sf ATI}_{{\mathcal{P}},D_{0}^{\Pi},\gamma}^{\epsilon,\delta}]-O(Q^{3}/s)-O(Q^{3}/\lvert{\mathcal{R}}\rvert)$
By taking ${\mathcal{R}}$ to be exponentially large and $Q$ is the number of queries the efficient ${\sf ATI}$ algorithm will make to the function $\Sigma,\Pi$ (in our case, $Q=O(1/\epsilon^{2}\cdot 1/\log\delta)\cdot T\cdot{\sf poly}(\lambda)$ for some polynomial $T$, where T is the running time of applying the controlled projection Definition 5.6 in our setting once).
Hybrid 3
[…]
Hybrid 4
Now we use the distribution $D_{0}$, with everything else staying the same. That is, we use distribution $D_{1}^{G\circ E}$ on any input $r$ is $D_{1}(G(E(r)))$. We can view $D_{b}(G(\cdot))$ as a list of $s$ samples from $D_{b}$, which the input $r$ is mapped to an index $i\in[s]$ and $i$ selects which sample to use. Since ${\mathcal{D}}_{0}$ and ${\mathcal{D}}_{1}$ are statistically close by distance $\eta$, which requires $\Omega(1/\eta^{2})$ samples to distinguish. Then setting $s=1/(\eta^{\frac{3}{2}})$ samples will not help distinguish them. Therefore by combining hybrid 0-4 we have: $\operatorname{Tr}[{\sf ATI}_{{\mathcal{P}},D_{0},\gamma}^{\epsilon,\delta}]\geq\operatorname{Tr}[{\sf ATI}_{{\mathcal{P}},D_{1}^{G\circ E},\gamma}^{\epsilon,\delta}]-O(Q^{3}\cdot\eta^{\frac{3}{2}})-O(Q^{3}/\lvert{\mathcal{R}}\rvert)$
Hybrid 5-8
These hybrids will be the reverse of hybrid 0-4, with distribution ${\mathcal{D}}_{1}$, so we simply multiply the error terms by 2 and conclude the statement.
C.3 Proof for the Probability Relations in Section 10.3
@@ Appendix D Quantum Search-to-Decision LWE with Almost-Perfect Extraction(Clean Version)
Let $\lambda\in\mathbb{N}$ be the security parameter. Let $n,m\in\mathbb{N}$ be integers and let $q$ be a prime. If there is a $\gamma$-good quantum distinguisher for decisional ${\sf LWE}_{n,m,q,\sigma}$ with some auxiliary quantum input running in time $T$ for some noticeable $\gamma$, there there exists a quantum algorithm that solves search ${\sf LWE}_{n,m,q,\sigma}$ with probability $(1-\mathsf{negl}(\lambda))$, running in time $T^{\prime}=Tnp\cdot\mathsf{poly}(1/\epsilon,1/\log\delta)$ where $p$ is the size of the domain for each entry $\mathbf{s}_{i}$ in the secret $\mathbf{s}$.  [^11] Our ${\sf LWE}_{n,m,q,\sigma}$ instance may be over a large modulus $q$, but we require the secret $\mathbf{s}$ to have entries in some smaller domain $[k,k+p)\in\mathbb{Z}_{q}$ of size $p$. Our algorithm will work for polynomial or superpolynomial but subexponential $p$. and $\epsilon,\delta\in(0,1)$ [^12] In our algorithm, we will set $\epsilon$ to $\gamma/100pn$ and set $\delta$ to be exponentially small..

### 06 Mar 2024 v4

@@ 10.2.1 Proof Outline Using Inefficient Measurement ${\sf TI}$
To make the above algorithm efficient, we use an approximate version of ${\sf TI}$, called ${\sf ATI}$ (seeSection 5.2 for detailed prelims.). When we move on to using an approximate version of ${\sf TI}$, we will have to be more carefull: negligible errors will accumulate additively on the measurement outcomes $\rho_{\mathsf{Delete}}$ after every ${\sf TI}$-measurement, but will stay negligible throughout the algorithm with our choice of parameters (see Section 11 for details) by the properties of ATI in Section 5.2. By these guarantees, we obtain $\mathbf{x}$ with $1$ probability in the end.
@@ 11.2 Extraction Algorithm
  • Let $\mathbf{x}_{\ell,j}^{\prime}$ be the register that stores the final guess for the $j$-th entry of $\mathbf{x}_{\ell}$, initialized with all zeros.
@@
          Run ${\sf ATI}_{1/2+\gamma,{\mathcal{P}},{\mathcal{D}}(g_{\ell,j})}^{\epsilon,\delta}$ on $\rho_{\mathsf{Delete}}$ wih respect to $\mathsf{pk}$.
@@ 11.3 Analysis of the Extractor
  • If ${\sf ATI}^{\epsilon,\delta}_{{\mathcal{P}},{\mathcal{D}}_{2.k},1/2+\gamma}\,\rho$ outputs 1 and a leftover state $\rho^{\prime}$, then $\operatorname{Tr}[{\sf ATI}^{\epsilon,\delta}_{{\mathcal{P}},{\mathcal{D}}(g_{\ell,j})),1/2+\gamma-3\epsilon}\,\rho^{\prime}]\geq 1-2\delta-\mathsf{negl}(\lambda)$.
@@
This follows directly from 11.1, Lemma 5.8 and Corollary 5.10.
By Lemma 5.8, we have that if ${\sf ATI}^{\epsilon,\delta}_{{\mathcal{P}},{\mathcal{D}}_{2.k},1/2+\gamma}\,\rho$ outputs 1 and a leftover state $\rho^{\prime}$, then $\ \operatorname{Tr}[{\sf ATI}^{\epsilon,\delta}_{{\mathcal{P}},{\mathcal{D}}_{2.k},1/2+\gamma-3\epsilon}\,\rho^{\prime}]\geq 1-\delta$. Next, by Corollary 5.10, we have $\operatorname{Tr}[{\sf ATI}^{\epsilon,\delta}_{{\mathcal{P}},{\mathcal{D}}(g_{\ell,j})),1/2+\gamma}\,\rho^{\prime}]\geq\operatorname{Tr}[{\sf ATI}^{\epsilon,\delta}_{{\mathcal{P}},{\mathcal{D}}_{2.k},1/2+\gamma}\,\rho^{\prime}]-\eta_{0}^{\prime}$. Thus, overall we have $\operatorname{Tr}[{\sf ATI}^{\epsilon,\delta}_{{\mathcal{P}},{\mathcal{D}}(g_{\ell,j})),1/2+\gamma-3\epsilon}\,\rho^{\prime}]\geq 1-\delta-\mathsf{negl}(\lambda)=1-\mathsf{negl}(\lambda)$ since $\delta$ is exponentially small. ∎
@@
By Corollary 11.3, have that $\operatorname{Tr}[{\sf ATI}^{\epsilon,\delta}_{{\mathcal{P}},D(g_{i}),1/2+\gamma-4\epsilon}\,\rho]\geq 1-3\delta-\mathsf{negl}(\lambda)\geq 1-\mathsf{negl}(\lambda)$, where $\delta$ is exponentially small.
@@
When the $L+1$-th measurement ${\sf ATI}^{\epsilon,\delta}_{{\mathcal{P}},D(g_{\ell,j}),1/2+\gamma}$ uses a correct $g_{\ell,j}$, then with probability $(1-L^{\prime}\cdot O(\eta_{1}\cdot\ln(4/\delta)/\epsilon))(1-\mathsf{negl}(\lambda))$, we obtain outcome $1$, where $L^{\prime}$ is the number of times we get outcome 0 since the last time we get outcome 1, by the fact that $|\operatorname{Tr}({\mathcal{P}}\rho_{L-L^{\prime}})-\operatorname{Tr}({\mathcal{P}}\rho_{L})|\leq\lVert\rho_{L-L^{\prime}}-\rho_{L}\rVert_{\operatorname{Tr}}$ for all POVM measurements ${\mathcal{P}}$.
The total probability we extract all $k\cdot n$-entries of the LWE secret is: $(1-\mathsf{negl}(\lambda))^{kn}\cdot(1-\delta)^{O(B_{X}\cdot kn)}\cdot(1-O(\eta_{1}\cdot\ln(4/\delta)/\epsilon))^{O(B_{X}\cdot kn)}$. Here, $(1-\mathsf{negl}(\lambda))^{kn}$ is the total loss incurred by measurements ${\sf ATI}^{\epsilon,\delta}_{{\mathcal{P}},D(g_{\ell,j}),1/2+\gamma}$ with correct $g_{\ell,j}$’s, because we will make $kn$ (the dimension of the secret) number of such measurements; the error $(1-\delta)^{O(B_{X}\cdot kn)}\cdot(1-O(\eta_{1}\cdot\ln(4/\delta)/\epsilon))^{O(B_{X}\cdot kn)}$ is incurred by measurements ${\sf ATI}^{\epsilon,\delta}_{{\mathcal{P}},D(g_{\ell,j}),1/2+\gamma}$ with incorrect $g_{\ell,j}$’s, because we will make $O(B_{X}kn)$ (the dimension of the secret multiplied by the range of the secret) number of such measurements.
@@
Let us recall our parameter settings (the followings are all in terms of security parameter $\lambda$): $q^{-n}$ is exponentially small and thus $\eta_{1}$ is exponentially small; $\delta$ is exponentially small. $k,n$ are polynomial. $B_{X}$ is quasipolynomial.
@@
$(1-\mathsf{negl}(\lambda))^{kn}\cdot(1-\delta)^{O(B_{X}\cdot kn)}\cdot(1-O(\eta_{1}\cdot\ln(4/\delta)/\epsilon))^{O(B_{X}\cdot kn)}$
$\geq$ $(1-kn\cdot\mathsf{negl}(\lambda))\cdot(1-O(B_{X}kn\cdot\delta))\cdot(1-O(B_{X}kn\cdot\eta_{1}\cdot\ln(4/\delta)/\epsilon))$
$\geq$ $1-kn\cdot\mathsf{negl}(\lambda)$
@@ Appendix C Additional Missing Proofs
C.2 Proof for the Probability Relations in Section 10.3
@@ Appendix D Quantum Search-to-Decision LWE with Almost-Perfect Extraction(Clean Version)
Let $\lambda\in\mathbb{N}$ be the security parameter. Let $n,m\in\mathbb{N}$ be integers and let $q$ be a prime. If there is a $\gamma$-good quantum distinguisher for decisional ${\sf LWE}_{n,m,q,\sigma}$ with some auxiliary quantum input running in time $T$ for some noticeable $\gamma$, there there exists a quantum algorithm that solves search ${\sf LWE}_{n,m,q,\sigma}$ with probability $(1-\mathsf{negl}(\lambda))$, running in time $T^{\prime}=Tnp\cdot\mathsf{poly}(1/\epsilon,1/\log\delta)$ where $p$ is the size of the domain for each entry $\mathbf{s}_{i}$ in the secret $\mathbf{s}$. and $\epsilon,\delta\in(0,1)$ [^12] In our algorithm, we will set $\epsilon$ to $\gamma/100pn$ and set $\delta$ to be exponentially small..